News

Las Cumbres Observatory Provides Vital Early Observations of a Supernova

Earth and Space workshop 2017

Last week (Friday 3 February) the LCO education team hosted a teacher training event in Cardiff University, jointly with School of Physics & Astronomy, Universe in the Classroom, and Space Awareness. It was the not only the largest teacher training event ever run at the School of Physics & Astronomy (with 35 registered teachers from across Wales) but was also our first event as official partners with Space Awareness.

New Key Projects announced

Following the recommendations of an independent Time Allocation Committee (TAC), five new Key Projects have been awarded time on the LCO Network. The list of all key projects that will run during the 2017AB semester has been posted on the website's Science section. 

Call for proposals 2017AB

Las Cumbres Observatory is soliciting proposals for science observations for the 2017AB semester, which will begin on 1 April 2017 and run through 30 November 2017. This call is for astronomers from institutions without guarantees of Network time: the University of California at Santa Barbara (UCSB), the Infrared Processing and Analysis Center (IPAC), the University of Colorado, and the Instituto de Astrofisica de Canarias (IAC). Astronomers affiliated with LCO, including Time Allocation Committee members and Science Advisory Committee members, are also welcome to submit proposals. (LCO's Science Collaboration partners that have guaranteed time conduct independent proposal solicitations and reviews.)

Network status: Happy Holidays!

The Network suffered a one-two punch of power outages in the past week. On December 17, power was lost on Haleakala. As a result, science observations from that site (OGG) weren't possible on December 17-18. ----- A few days earlier, on December 13, there were power outages that affected all CTIO facilities. Consequently, we made no science observation from our LSC site on December 13-14. Since then, the cameras in Dome B (fl03) and Dome C (fl04) at LSC have been showing unstable temperatures. As a safeguard, we will warm these cameras up tonight (2016-12-21), then inspect and recharge their cryogenic systems tomorrow. Thus, only one camera (fl15, in Dome A) will be available for science observations tonight. We expect to have all cameras available again for the night of 2016-12-22.